OpenStudy (anonymous):

Simplify and write with positive exponents: (x^3-125)/(x^2-25)

OpenStudy (curry):

x+5 x-5 x+5/x+5 x-5=x+5

jimthompson5910 (jim_thompson5910):

(x^3-125)/(x^2-25) ((x-5)(x^2+5x+25)/(x^2-25) ((x-5)(x^2+5x+25)/((x-5)(x+5)) (x^2+5x+25)/(x+5) So \[\large \frac{x^3-125}{x^2-25}=\frac{x^2+5x+25}{x+5}\]

OpenStudy (curry):

oh wait i messed up good catch jim

OpenStudy (anonymous):

Thanks Jim!

jimthompson5910 (jim_thompson5910):

you factor the numerator using the difference of cubes formula you factor the denominator using the difference of squares formula